OpenStudy (anonymous):

Use rules of logarithms to expand. log (x+3)^2/(x-2)(x^2+5)^4

OpenStudy (solomonzelman):

\[Log(~~(x+3)^2/(x-2)(x^2+5)^4~~)\]

OpenStudy (solomonzelman):

Do you know the rules? Any ideas?

OpenStudy (anonymous):

No. I'm still trying to learn this stuff. It's a little hard for me to understand.

OpenStudy (solomonzelman):

\[is~~(x^2+5)^4~~part ~~of~~the~~denominator?\]

OpenStudy (solomonzelman):

I think yes, so use\[\log(a/b)=\log(a)-\log(b)\]

OpenStudy (anonymous):

It is a part of the denominator.

OpenStudy (solomonzelman):

\[2\log(x+3)-\log(x-2)+4\log(x^2+5)\]

OpenStudy (solomonzelman):

there

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

Anytime!

OpenStudy (solomonzelman):

actually. \[2\log(x+3)-log(x-2)\color{red}{-}4\log(x^2+5)\]

OpenStudy (solomonzelman):

my bad, for the mistake, this one is right I think.

OpenStudy (anonymous):

ok